# Shared Global and Local Geometry of Language Model Embeddings

### v1

@@ Abstract
Researchers have recently suggested that models share common representations. In this work, we find that the token embeddings of language models exhibit common geometric structure. First, we find “global” similarities: token embeddings often share similar relative orientations. Next, we characterize local geometry in two ways: (1) by using Locally Linear Embeddings, and (2) by defining a simple measure for the intrinsic dimension of each token embedding. Our intrinsic dimension measure demonstrates that token embeddings lie on a lower dimensional manifold. We qualitatively show that tokens with lower intrinsic dimensions often have semantically coherent clusters, while those with higher intrinsic dimensions do not. Both characterizations allow us to find similarities in the local geometry of token embeddings. Perhaps most surprisingly, we find that alignment in token embeddings persists through the hidden states of language models, allowing us to develop an application for interpretability. Namely, we empirically demonstrate that steering vectors from one language model can be transferred to another, despite the two models having different dimensions.
@@ 1 Introduction
Perhaps most surprising is that the alignment in token embeddings seem to persist in the hidden layers of the language models. Given these insights, we demonstrate an application for model interpretability: steering vectors that can control one model can be transferred and reused for another model, despite the models having different dimensions.
@@ 5.4 Similar Intrinsic Dimensions Across Language Models
6 Application: Transferring Steering Vectors
Here we build on our insights to demonstrate an application for model interpretability. Namely, we show that steering vectors (Section 2) can be transferred across language models.
@@ 6 Application: Transferring Steering Vectors
Our approach is simple. Given “source” and “target” models $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$, we randomly sample a set of $N$ (= 100,000) tokens. We notate the $N$ unembedding vectors from the two models as $\mathcal{U}_{S}$ and $\mathcal{U}_{T}$. We fit a linear transformation, $A$, to map points $\mathcal{U}_{S}$ to $\mathcal{U}_{T}$, using least squares minimization. Note that $A$ maps between spaces with different dimensions.
@@ 7 Discussion
Perhaps most surprisingly, we demonstrate that alignment in the model’s embedding space can persist through the hidden states of the model. This leads to an application for model interpretability: steering vectors can be transferred across language models.

### v2

@@ Abstract
Researchers have recently suggested that models share common representations. In our work, we find that token embeddings of language models exhibit common geometric structure. First, we find “global” similarities: token embeddings often share similar relative orientations. Next, we characterize local geometry in two ways: (1) by using Locally Linear Embeddings, and (2) by defining a simple measure for the intrinsic dimension of each token embedding. Our intrinsic dimension demonstrates that token embeddings lie on a lower dimensional manifold. We qualitatively show that tokens with lower intrinsic dimensions often have semantically coherent clusters, while those with higher intrinsic dimensions do not. Both characterizations allow us to find similarities in the local geometry of token embeddings. Perhaps most surprisingly, we find that alignment in token embeddings persists through the hidden states of language models, allowing us to develop an application for interpretability. Namely, we introduce Emb2Emb, a simple method to transfer steering vectors from one language model to another, despite the two models having different dimensions.
@@ 1 Introduction
Perhaps most surprising is that the alignment in token embeddings seem to persist in the hidden layers of the language models. Given these insights, we introduce Emb2Emb, a simple tool for model interpretability: steering vectors that can control one model can be transferred and reused for another model, despite the models having different dimensions.
@@ 5.4 Similar Intrinsic Dimensions Across Language Models
6 Emb2Emb: Transferring Steering Vectors
Based on our insights, we introduce Emb2Emb, a simple tool for model interpretability. Namely, we show that steering vectors (Section 2) can be transferred across language models.
@@ 6 Emb2Emb: Transferring Steering Vectors
Emb2Emb is simple. Given “source” and “target” models $\mathcal{M}_{S}$ and $\mathcal{M}_{T}$, we randomly sample a set of $N$ (= 100,000) tokens. We notate the $N$ unembedding vectors from the two models as $\mathcal{U}_{S}$ and $\mathcal{U}_{T}$. We fit a linear transformation, $A$, to map points $\mathcal{U}_{S}$ to $\mathcal{U}_{T}$, using least squares minimization. Note that $A$ maps between spaces with different dimensions.
@@ 7 Discussion
Perhaps most surprisingly, alignment in the model’s embedding space seem to persist through the hidden states. This leads to Emb2Emb, a simple application for model interpretability: steering vectors can be transferred across language models.